OpenStudy (anonymous):

i really need help please.Gino designed a garden for his rectangular backyard. The design he drew on paper is 10 inches long. The length of his backyard is 40 feet. If his drawing is to scale, and his backyard is 28 feet wide, how many inches wide is his drawing? Write a proportion using words. Then, use the proportion to calculate the drawing width. Show your work. i dont need the answer i just need the porportion with words please.

OpenStudy (anonymous):

10/40 = x/28 x = 7 7 Inches wide

OpenStudy (anonymous):

thanks but can you tell me how to right the porportion with words